Phoebe,Eloise and Radhika have the combined age of 88. Eloise is 3 years older than Radhika, who is twice as old as Phoebe, how old is Eloise?
Answer by Cromlix(4381)

Hi there,
Eloise is 3 years older than Radhika,
who is twice as old as Phoebe,
Make Phoebe's age = x
Radhika's age then is = 2x
Eloise's age = 2x + 3
Phoebe,Eloise and Radhika have
the combined age of 88.
x + 2x + 2x + 3 = 88
Collect like terms:
x + 2x + 2x = 88 - 3
5x = 85
x = 17
Phoebe is 17 years old
Radhika is 34 years old
Eloise is 37 years old.
Hope this helps :-)